the price of a product was decreased by 20% and the reduced price still by 10%. How many percent was the total reduction of the original price?
Answer by Theo(13342)

p * (1-.20) * (1-.10) = p * .80 * .90 = p * .72

the total reduction in price was p - .72 * p = .28 * p = 28% of p.

if the price was 100, then:
100 - .20 * 100 = 100 - 20 = 80
80 - .10 * 80 = 80 - 8 = 72
the final sale price is 72.
the reduction in price is 100 - 72 = 28.
the reduction percent is 28/100 * 100 = 28%.